Suppose you have two charges q1 and q2, and both are located on the x-axis. You know
that q1 has a charge of 7.58
1
Expert's answer
2020-05-08T16:12:48-0400
"F=\frac{kq_1q_2}{r^2}=7.58\frac{kq_1}{r^2}"